# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 2.2. Linear and nonlinear operators
$\mathscr{A}\boldsymbol{u}-\mathscr{A}\boldsymbol{v},\boldsymbol{w}\right>$ $=(\nabla(\boldsymbol{u}-\boldsymbol{v}),\nabla\boldsymbol{w})$
@@
Hence, the operator $\mathscr{K}(\cdot):\mathbb{X}\to\mathbb{X}^{\prime}$ is locally Lipschitz. ∎
2.3. Stochastic third-grade fluids equations
@@
[…]
where $\boldsymbol{x}\in\mathbb{H},\ \boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a two-sided cylindrical Wiener process in $\mathbb{H}$ with its RKHS $\mathrm{K}$ which satisfies Hypothesis 1.1. Now we present the definition of weak solution (in the analytic sense) to system (2.18) with the initial data $\boldsymbol{x}\in\mathbb{H}$ at the initial time $s\in\mathbb{R}.$
@@
[…]
@@
3. RDS generated by stochastic third-grade fluids equations
In this section, we prove the existence of RDS generated by 3D stochastic third-grade fluids equations using a random transformation (known as Doss-Sussman transformation, [29, 59]). The process of generation of RDS has been adopted from the works [12, 13].
@@ 3.1. Metric dynamical system
$C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})$ =$=\left\{\omega\in C(\mathbb{R};\mathrm{E}):\omega(0)=\boldsymbol{0},\ \|\omega\|_{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}<\infty\right\},$
$\Omega(\xi,\mathrm{E})$ =$=\overline{\{\omega\in C^{\infty}_{0}(\mathbb{R};\mathrm{E}):\omega(0)=0\}}^{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}.$
@@
is an $\mathrm{E}$-valued two-sided Wiener process such that the RKHS of the Gaussian measure $\mathscr{L}(w_{1})$ on $\mathrm{E}$ is $\mathrm{K}$. For $t\in\mathbb{R},$ let $\mathscr{F}_{t}:=\sigma\{w_{s}:s\leq t\}.$ Then there exists a unique bounded linear map $\mathrm{W}(t):\mathrm{K}\to\mathrm{L}^{2}(\Omega(\xi,\mathrm{E}),\mathscr{F}_{t},\mathbb{P}).$ Moreover, the family $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a $\mathrm{K}$-cylindrical Wiener process on a filtered probability space $(\Omega(\xi,\mathrm{E}),\mathscr{F},\{\mathscr{F}_{t}\}_{t\in\mathbb{R}},\mathbb{P})$ (cf. [14] for more details).
@@ 3.2. Ornstein-Uhlenbeck process
Proposition 3.2 ([12, Proposition 2.11]).
Let $-\mathbb{A}$ be the generator of an analytic semigroup $\{e^{t\mathbb{A}}\}_{t\geq 0}$ on a separable Banach space $\mathbb{Y}$ such that for some $C>0\ \text{and}\ \gamma>0$
@@
$\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\ni\widetilde{\omega}\mapsto\hat{\boldsymbol{z}}(t)\in\mathbb{Y}$
is continuous. Moreover, the map $\hat{\boldsymbol{z}}:\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\to\mathrm{C}_{1/2}(\mathbb{R};\mathbb{Y})$ is well defined, linear and bounded. In particular, there exists a constant $C>0$ such that for any $\widetilde{\omega}\in\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})$
@@
    For all $-\infty<a<b<\infty$ and $t\in\mathbb{R}$, the map
    $\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\ni\widetilde{\omega}\mapsto(\hat{\boldsymbol{z}}(\widetilde{\omega})(t),\hat{\boldsymbol{z}}(\widetilde{\omega}))\in\mathbb{Y}\times\mathrm{L}^{q}(a,b;\mathbb{Y}),$ (3.5)
@@
Moreover, $\mathbb{E}\left[\|\boldsymbol{z}_{\chi}(t)\|^{2}_{\mathfrak{X}}\right]\to 0$ as $\chi\to\infty.$
@@
[…]
@@
$\lim_{t\to-\infty}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ e^{ct}=0.$
@@
Let us fix $\omega\in\Omega$. Because of (3.4), there exists a $\rho_{1}=\rho_{1}(\omega)\geq 0$ such that
@@
$\lim_{t\to-\infty}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ e^{ct}\leq$ $\rho_{1}^{2}\lim_{t\to-\infty}|t|^{2}e^{ct}=0,$
@@
A function $\kappa:\Omega\to(0,\infty)$ belongs to the class $\mathfrak{K}$ if and only if
$\lim_{t\to\infty}[\kappa(\theta_{-t}\omega)]^{2}e^{-ct}=0,$ (3.16)
@@
Let us denote the class of all closed and bounded random sets $D$ on $\mathbb{H}$ by $\mathfrak{DK}$ such that the radius function $\Omega\ni\omega\mapsto\kappa(D(\omega)):=\sup\{\|x\|_{2}:x\in D(\omega)\}$ belongs to the class $\mathfrak{K}.$ It is straight forward that the constant functions belong to $\mathfrak{K}$. It is clear by Definition 3.1 that the class $\mathfrak{K}$ is closed with respect to sum, multiplication by a constant and if $\kappa\in\mathfrak{K},0\leq\bar{\kappa}\leq\kappa,$ then $\bar{\kappa}\in\mathfrak{K}.$
@@
We define functions $\kappa_{i}:\Omega\to(0,\infty),i=1,\ldots,4,$ by the following formulae: for $\omega\in\Omega,$
$[\kappa_{1}(\omega)]^{2}$ $:=\|\boldsymbol{z}_{\chi}(\omega)(0)\|_{2},$ $[\kappa_{2}(\omega)]^{2}:=\sup_{s\leq 0}\|\boldsymbol{z}_{\chi}(\omega)(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s},$
$[\kappa_{3}(\omega)]^{2}$ $:=\int_{-\infty}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ \mathrm{d}t,$ $[\kappa_{4}(\omega)]^{2}:=\int_{-\infty}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{4}_{\mathbb{W}^{1,4}}\ \mathrm{d}t.$
@@
$\lim_{t\to\infty}[\kappa_{1}(\theta_{-t}\omega)]^{2}e^{-ct}=$ $\lim_{t\to\infty}\|\boldsymbol{z}_{\chi}(\theta_{-t}\omega)(0)\|^{2}_{2}e^{-ct}=\lim_{t\to\infty}\|\boldsymbol{z}_{\chi}(\omega)(-t)\|^{2}_{2}e^{-ct}.$
@@
$\lim_{t\to\infty}[\kappa_{2}(\theta_{-t}\omega)]^{2}e^{-ct}\leq\lim_{t\to\infty}e^{-(c-c_{1})t}\lim_{t\to\infty}\left[\sup_{\sigma\leq-t}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}\ e^{c_{1}\sigma}\right]=0,$
@@
$\to 0\ \text{ as }\ t\to\infty.$

### v2

@@ 2.3. Linear and nonlinear operators
$\left<\mathscr{A}\boldsymbol{u}-\mathscr{A}\boldsymbol{v},\boldsymbol{w}\right$> $=(\nabla(\boldsymbol{u}-\boldsymbol{v}),\nabla\boldsymbol{w})$
@@
Hence, the operator $\mathscr{G}(\cdot):\mathbb{X}\rightarrow\mathbb{X}^{\prime}$ is locally Lipschitz. ∎
2.4. A compact operator
We follow [14, Subsection 2.3] for the contents of this section. Consider the natural embedding $j:\mathbb{V}\hookrightarrow\mathbb{H}$ and its adjoint $j^{*}:\mathbb{H}\hookrightarrow\mathbb{V}$. Since the range of $j$ is dense in $\mathbb{H}$, the map $j^{*}$ is one-to-one. Let us define
$\mathrm{D}(\mathbb{A})$ $:=j^{*}(\mathbb{H})\subset\mathbb{V},$
$\mathbb{A}\boldsymbol{v}$ $:=(j^{*})^{-1}\boldsymbol{v},\ \ \boldsymbol{v}\in\mathrm{D}(\mathbb{A}).$ (2.24)
Note that for all $\boldsymbol{v}\in\mathrm{D}(\mathbb{A})$ and $\boldsymbol{v}\in\mathbb{V}$
$(\mathbb{A}\boldsymbol{u},\boldsymbol{v})_{\mathbb{H}}=(\boldsymbol{u},\boldsymbol{v})_{\mathbb{V}}.$
Let us assume that $s>2$. It is clear that $\mathbb{V}_{s}$ is dense in $\mathbb{V}$ and the embedding $j_{s}:\mathbb{V}_{s}\hookrightarrow\mathbb{V}$ is continuous. Then, there exists a Hilbert space $\mathbb{U}$ (cf. [39], [14, Lemma C.1]) such that $\mathbb{U}\subset\mathbb{V}_{s}$, $\mathbb{U}$ is dense in $\mathbb{V}_{s}$ and
$\text{ the natural embedding }\iota_{s}:\mathbb{U}\hookrightarrow\mathbb{V}_{s}\text{ is compact.}$
It implies that
[…]
Consider the composition
$\iota:=j\circ j_{s}\circ\iota_{s}:\mathbb{U}\rightarrow\mathbb{H}$
and its adjoint
$\iota^{*}:=(j\circ j_{s}\circ\iota_{s})^{*}=\iota_{s}^{*}\circ j^{*}_{s}\circ j^{*}:\mathbb{H}\rightarrow\mathbb{U}.$
We have that $\iota$ is compact and since its range is dense in $\mathbb{H}$, $\iota^{*}:\mathbb{H}\rightarrow\mathbb{U}$ is one-one. Let us define
$\mathrm{D}(\mathscr{L})$ $:=\iota^{*}(\mathbb{H})\subset\mathbb{U},$
$\mathscr{L}\boldsymbol{v}$ $:=(\iota^{*})^{-1}\boldsymbol{v},\ \ \boldsymbol{v}\in\mathrm{D}(\mathscr{L}).$ (2.25)
Also we have that $\mathscr{L}:\mathrm{D}(\mathscr{L})\rightarrow\mathbb{H}$ is onto, $\mathrm{D}(\mathscr{L})$ is dense in $\mathbb{H}$ and
$(\mathscr{L}\boldsymbol{v},\boldsymbol{w})_{\mathbb{H}}=(\boldsymbol{v},\boldsymbol{w})_{\mathbb{U}},\ \ \ \ \boldsymbol{v}\in\mathrm{D}(\mathscr{L}),\ \boldsymbol{w}\in\mathbb{U}.$
Furthermore, for $\boldsymbol{v}\in\mathrm{D}(\mathscr{L})$,
$\mathscr{L}\boldsymbol{v}=(\iota^{*})^{-1}\boldsymbol{v}=(j^{*})^{-1}\circ(j^{*}_{s})^{-1}\circ(\iota^{*}_{s})^{-1}\boldsymbol{v}=\mathbb{A}\circ(j^{*}_{s})^{-1}\circ(\iota^{*}_{s})^{-1}\boldsymbol{v},$
where $\mathbb{A}$ is defined in (2.4). Since the operator $\mathscr{L}$ is self-adjoint and $\mathscr{L}^{-1}$ is compact, there exists an orthonormal basis $\{\boldsymbol{e}_{i}\}_{i\in\mathbb{N}}$ of $\mathbb{H}$ such that
$\mathscr{L}\boldsymbol{e}_{i}=\mu_{i}\boldsymbol{e}_{i},\ \ \ \ i\in\mathbb{N},$ (2.26)
that is, $\{\boldsymbol{e}_{i}\}$ are the eigenfunctions and $\{\mu_{i}\}$ are the corresponding eigenvalues of operator $\mathscr{L}$. Note that $\boldsymbol{e}_{i}\in\mathbb{U}$, $i\in\mathbb{N}$, because $\mathrm{D}(\mathscr{L})\subset\mathbb{U}$.
Let us fix $m\in\mathbb{N}$ and let $\mathrm{P}_{m}$ be the operator from $\mathbb{U}^{\prime}$ to $\mathrm{span}\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}\}=:\mathbb{H}_{m}$ defined by
$\mathrm{P}_{m}\boldsymbol{v}^{*}:=\sum_{i=1}^{m}\langle\boldsymbol{v}^{*},\boldsymbol{e}_{i}\rangle_{\mathbb{U}^{\prime}\times\mathbb{U}}\boldsymbol{e}_{i},\ \ \ \ \ \ \boldsymbol{v}^{*}\in\mathbb{U}^{\prime}.$ (2.27)
We will consider the restriction of operator $\mathrm{P}_{m}$ to the space $\mathbb{H}$ denoted still by the same. In particular, we have $\mathbb{H}\hookrightarrow\mathbb{U}^{\prime}$, that is, every element $\boldsymbol{v}\in\mathbb{H}$ induces a functional $\boldsymbol{v}^{*}\in\mathbb{U}^{\prime}$ by
$\langle\boldsymbol{v}^{*},\boldsymbol{u}\rangle_{\mathbb{U}^{\prime}\times\mathbb{U}}:=(\boldsymbol{v},\boldsymbol{u}),\ \ \ \ \boldsymbol{u}\in\mathbb{U}.$ (2.28)
Thus the restriction of $\mathrm{P}_{m}$ to $\mathbb{H}$ is given by
$\mathrm{P}_{m}\boldsymbol{v}:=\sum_{i=1}^{m}(\boldsymbol{v},\boldsymbol{e}_{i})\boldsymbol{e}_{i},\ \ \ \ \ \ \boldsymbol{v}\in\mathbb{H}.$ (2.29)
Hence, in particular, $\mathrm{P}_{m}$ is the orthogonal projection from $\mathbb{H}$ onto $\mathbb{H}_{m}$.
Lemma 2.6 (See [14, Lemma 2.4]).
For every $\boldsymbol{v}\in\mathbb{U}$ and $s>2$, we have
  • (i)
    $\|\mathrm{P}_{m}\boldsymbol{v}\|_{\mathbb{U}}\leq\|\boldsymbol{v}\|_{\mathbb{U}}$,
  • (ii)
    $\lim\limits_{m\rightarrow\infty}\|\mathrm{P}_{m}\boldsymbol{v}-\boldsymbol{v}\|_{\mathbb{U}}=0$,
  • (iii)
    $\lim\limits_{m\rightarrow\infty}\|\mathrm{P}_{m}\boldsymbol{v}-\boldsymbol{v}\|_{\mathbb{V}_{s}}=0$,
  • (iv)
    $\lim\limits_{m\rightarrow\infty}\|\mathrm{P}_{m}\boldsymbol{v}-\boldsymbol{v}\|_{\mathbb{V}}=0$,
  • (v)
    $\lim\limits_{m\rightarrow\infty}\|\mathrm{P}_{m}\boldsymbol{v}-\boldsymbol{v}\|_{\mathbb{W}^{1,4}}=0$.
2.5. A subclass of stochastic third-grade fluids equations
@@
$\left\{\begin{aligned} \mathrm{d}\boldsymbol{v}(t)+\{\nu\mathscr{A}\boldsymbol{v}(t)+\mathscr{B}(\boldsymbol{v}(t))+\alpha\mathscr{J}(\boldsymbol{v}(t))+\beta\mathscr{K}(\boldsymbol{v}(t))\}\mathrm{d}t&=\mathscr{P}\boldsymbol{f}\mathrm{d}t+\mathrm{d}\mathrm{W}(t),\ \ \ t>0,\\ \boldsymbol{v}(0)&=\boldsymbol{x},\end{aligned}\right$. (2.30)
where $\boldsymbol{x}\in\mathbb{H},\ \boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a two-sided cylindrical Wiener process in $\mathbb{H}$ with its RKHS $\mathrm{K}$ which satisfies Hypothesis 1.1. Now we present the definition of weak solution (in the analytic sense) to system (2.30) with the initial data $\boldsymbol{x}\in\mathbb{H}$ at the initial time $s\in\mathbb{R}.$
@@
Suppose that (1.4) and Hypotheses 1.1-1.2 are satisfied. If $\boldsymbol{x}\in\mathbb{H}$, $s\in\mathbb{R}$, $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a two-sided Wiener process with its RKHS $\mathrm{K}$. A process $\{\boldsymbol{v}(t),\ t\geq s\},$ with trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ is a solution to system (2.30) if and only if $\boldsymbol{v}(s)=\boldsymbol{x}$ and for any $\boldsymbol{\phi}\in\mathbb{X}$, $t>s,$ $\mathbb{P}$-a.s.,
@@
3. RDS generated by system (1.5)
In this section, we prove the existence of RDS generated by system (1.5) using a random transformation (known as Doss-Sussmann transformation, [30, 67]). The process of generation of RDS has been adopted from the works [13, 10].
@@ 3.1. Metric dynamical system
$C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})$ $=\left\{\omega\in C(\mathbb{R};\mathrm{E}):\omega(0)=\boldsymbol{0},\ \|\omega\|_{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}<\infty\right\},$
$\Omega(\xi,\mathrm{E})$ $=\overline{\{\omega\in C^{\infty}_{0}(\mathbb{R};\mathrm{E}):\omega(0)=0\}}^{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}.$
@@
is an $\mathrm{E}$-valued two-sided Wiener process such that the RKHS of the Gaussian measure $\mathscr{L}(w_{1})$ on $\mathrm{E}$ is $\mathrm{K}$. For $t\in\mathbb{R},$ let $\mathscr{F}_{t}:=\sigma\{w_{s}:s\leq t\}.$ Then there exists a unique bounded linear map $\mathrm{W}(t):\mathrm{K}\rightarrow\mathrm{L}^{2}(\Omega(\xi,\mathrm{E}),\mathscr{F}_{t},\mathbb{P}).$ Moreover, the family $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a $\mathrm{K}$-cylindrical Wiener process on a filtered probability space $(\Omega(\xi,\mathrm{E}),\mathscr{F},\{\mathscr{F}_{t}\}_{t\in\mathbb{R}},\mathbb{P})$ (cf. [15] for more details).
@@ 3.2. Ornstein-Uhlenbeck process
Proposition 3.2 ([13, Proposition 2.11]).
[…]
@@
$\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\ni\widetilde{\omega}\mapstochar\rightarrow\hat{\boldsymbol{z}}(t)\in\mathbb{Y}$
is continuous. Moreover, the map $\hat{\boldsymbol{z}}:\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\rightarrow\mathrm{C}_{1/2}(\mathbb{R};\mathbb{Y})$ is well defined, linear and bounded. In particular, there exists a constant $C>0$ such that for any $\widetilde{\omega}\in\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})$
@@
    For all $\infty<a<b<\infty$ and $t\in\mathbb{R}$, the map
    $\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\ni\widetilde{\omega}\mapstochar\rightarrow(\hat{\boldsymbol{z}}(\widetilde{\omega})(t),\hat{\boldsymbol{z}}(\widetilde{\omega}))\in\mathbb{Y}\times\mathrm{L}^{q}(a,b;\mathbb{Y}),$ (3.5)
@@
Moreover, $\mathbb{E}\left[\|\boldsymbol{z}_{\chi}(t)\|^{2}_{\mathfrak{X}}\right]\rightarrow 0$ as $\chi\rightarrow\infty.$
@@
$\left\{\begin{aligned} \mathrm{d}\boldsymbol{z}_{\chi}(t)+\left[(-\nu\Delta+\chi\mathrm{I})\boldsymbol{z}_{\chi}(t)+\nabla{p}\right]\mathrm{d}t&=\mathrm{d}\mathrm{W}(t),\ \ t\in\mathbb{R},\\ \mathrm{div}\;\boldsymbol{z}_{\chi}&=0,\end{aligned}\right$. (3.14)
@@
$\lim_{t\rightarrow-\infty}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ e^{ct}=0.$
@@
Let us fix $\omega\in\Omega$. Because of (3.4), there exist $\rho_{1}=\rho_{1}(\omega)\geq 0$ and $t_{0}\leq 0$ such that
@@
$\lim_{t\rightarrow-\infty}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ e^{ct}\leq$ $\rho_{1}^{2}\lim_{t\rightarrow-\infty}|t|^{2}e^{ct}=0,$
@@
A function $\kappa:\Omega\rightarrow(0,\infty)$ belongs to the class $\mathfrak{K}$ if and only if
$\lim_{t\rightarrow\infty}[\kappa(\theta_{-t}\omega)]^{2}e^{-ct}=0,$ (3.16)
@@
Let us denote the class of all closed and bounded random sets $D$ on $\mathbb{H}$ by $\mathfrak{DK}$ such that the radius function $\Omega\ni\omega\mapstochar\rightarrow\kappa(D(\omega)):=\sup\{\|x\|_{2}:x\in D(\omega)\}$ belongs to the class $\mathfrak{K}.$ It is straight forward that the constant functions belong to $\mathfrak{K}$. It is clear by Definition 3.1 that the class $\mathfrak{K}$ is closed with respect to sum, multiplication by a constant and if $\kappa\in\mathfrak{K},0\leq\bar{\kappa}\leq\kappa,$ then $\bar{\kappa}\in\mathfrak{K}.$
@@
We define functions $\kappa_{i}:\Omega\rightarrow(0,\infty),i=1,\ldots,4,$ by the following formulae: for $\omega\in\Omega,$
$[\kappa_{1}(\omega)]^{2}$ $:=\|\boldsymbol{z}_{\chi}(\omega)(0)\|_{2},\ \ \$$[\kappa_{2}(\omega)]^{2}:=\sup_{s\leq 0}\|\boldsymbol{z}_{\chi}(\omega)(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s},$
$[\kappa_{3}(\omega)]^{2}$ $:=\int_{-\infty}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ \mathrm{d}t,\ \ \$$[\kappa_{4}(\omega)]^{2}:=\int_{-\infty}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{4}_{\mathbb{W}^{1,4}}\ \mathrm{d}t.$
@@
$\lim_{t\rightarrow\infty}[\kappa_{1}(\theta_{-t}\omega)]^{2}e^{-ct}=$ $\lim_{t\rightarrow\infty}\|\boldsymbol{z}_{\chi}(\theta_{-t}\omega)(0)\|^{2}_{2}e^{-ct}=\lim_{t\rightarrow\infty}\|\boldsymbol{z}_{\chi}(\omega)(-t)\|^{2}_{2}e^{-ct}.$
@@
$\lim_{t\rightarrow\infty}[\kappa_{2}(\theta_{-t}\omega)]^{2}e^{-ct}\leq\lim_{t\rightarrow\infty}e^{-(c-c_{1})t}\lim_{t\rightarrow\infty}\left[\sup_{\sigma\leq-t}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}\ e^{c_{1}\sigma}\right]=0,$
@@
$\rightarrow 0\ \text{ as }\ t\rightarrow\infty.$